# Keeping in Place After the Storm—Emergency Assistance and Evictions

### v2

@@ 3.1.2 Hurricanes and Tropical Storms
Table 8: Baseline average monthly level per 1000 renters per ZIP code in the pre-event period.
Variable Control Group Treated Group
Eviction Rate 2.664 2.423
Filing Rate 3.950 3.829
@@
We expand our baseline results to include tropical storms to show that the increase in observed evictions and filings still holds when lower intensity storms are included. Although, the comparative magnitudes decreases, we show that this is largely due to tropical storms seeing far lower post-storm eviction rates. Hence, our result is likely to be robust from other temporal patterns that may have coincided in our analysis with the comparatively fewer hurricane events. Adding additional storm events ensures the increased evictions and filings results are likely due to coinciding with intense hurricane events and the additional absence of FEMA assistance.
@@
Figure 8 and Table 9 shows the baseline result even when including the greater number of tropical storm events still holds (note the three times increase in sample size). The decrease is largely driven by the inclusion of tropical storms, noting the treatment post-period coefficient drop from 0.371 to 0.128 for evictions when only tropical storms are included, although the decrease itself is not statistically significant.
@@
Figure 8: The figure provides the estimated coefficients from an event study examining the effect of storm exposure, including hurricanes and tropical storms on monthly eviction-related outcomes at the ZIP code level. The analysis covers a window from five months before to six months after the storm event, with the month immediately preceding the event (month -1) serving as the reference period. Panel (a) displays estimates for monthly eviction rates, while Panel (b) shows estimates for monthly filing rates. Vertical lines indicate 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
Figure 9: The figure provides the estimated coefficients from an event study examining the impact of tropical storm exposure on monthly eviction-related outcomes at the ZIP code level. The analysis spans a window from five months before to six months after the storm event, with the month immediately preceding the event (month -1) serving as the reference period. Panel (a) reports estimates for monthly eviction rates, while Panel (b) shows estimates for monthly filing rates. Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level. The results suggest no statistically significant change in either eviction or filing rates following a tropical storm event.
Table 9: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, combining both hurricanes and tropical storms. The analysis includes two specifications: one with all storms and another excluding hurricane events. The estimation window spans from 5 months before to 6 months after the event. In the full specification including hurricanes and tropical storms, the coefficient on Treated×$\times$×Post indicates a statistically significant increase of 0.209 evictions and 0.336 filings in treated ZIP codes. These correspond to approximately 8.6% and 8.7% increases relative to the treated group’s pre-event averages of 2.423 and 3.829, respectively. When hurricanes are excluded, the estimated effects are smaller and not statistically significant.
Dependent Variable Eviction Filing Eviction Filing
Period Monthly Monthly Monthly Monthly
Window [-5 ; 6] [-5 ; 6] [-5 ; 6] [-5 ; 6]
Treated×$\times$×Post 0.209∗∗∗ 0.336∗∗∗ 0.103 0.255
[0.070] [0.125] [0.105] [0.200]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Hurricanes Y Y N N
Observations 62209 62209 43708 43708
Table 10: Triple Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, combining both hurricanes and tropical storms. The analysis evaluates the differential effects of the tropical storm event across treated and control groups, with an estimation window spanning from 5 months before to 6 months after the event. The coefficient of 0.331 for evictions and 0.428 for filings in the Treated×$\times$×Post interaction term suggests a significant increase in both outcomes in the treated ZIP codes, corresponding to approximately 13.6% and 11.1% increases relative to the pre-event averages of 2.423 and 3.829, respectively. The triple interaction Treated×$\times$×Post×$\times$×Storm is negative and marginally significant (-0.228, significant at the 10%) for Eviction, suggesting that the treatment effect after tropical storms is somewhat lower than after hurricanes. In other words, tropical storms appear to have a weaker impact on evictions compared to hurricanes. However, for eviction filings, the triple interaction (-0.173) is negative but not statistically significant.
Dependent Variable Eviction Filing
Period Monthly Monthly
Window [-5 ; 6] [-5 ; 6]
Treated×$\times$×Post 0.331∗∗∗ 0.428∗∗∗
[0.089] [0.110]
Treated×$\times$×Post×$\times$×Storm -0.228 -0.173
[0.137] [0.212]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 62209 62209
@@ 3.2.1 Hurricanes
We further compare the usage of high cost credit during hurricane events with the availability of FEMA rental assistance. We find that both transaction volumes and default rates are lower post storm and that when FEMA assistance is made available, transactions volumes and default rates decrease by greater magnitude compared to when it is not made available, reflecting a reduction in demand of payday credit when FEMA is available but also a decrease in precarity in making repayments. In particular, Figure 10 we observe a decrease in defaults and reduction in transaction volumes (except a spike around 4,8 and 12 weeks, possibly reflecting a feature of payday loans that repeat customers often rollover their outstanding 30-day loans).
@@
Specifically, in Table 12 we show that the volume of transactions reduces by 8.364 per week and a decrease in defaults by 1.458 per week on average per ZIP code. This reflects a 3.8% decrease in transactions per week relative to the baseline period and 14% decrease relative to the baseline period. Comparatively, when FEMA rental assistance is available, we find transactions volumes decrease by 12.597 per week relative to non-FEMA ZIP codes, a decline of 5.7% relative to the pre-event baseline. Furthermore, we find a decrease decrease in defaults by 3.318 per week, an approximately 31.8% relative to the pre-event baseline. While when ZIP codes receiving FEMA assistance are excluded, transaction volumes reduced by 8.538 per ZIP code per week relative to the baseline period, a decrease of 3.9% per week per ZIP code relative to the baseline period. Defaults, on the other hand, show a not statistically significant decrease of 0.433.
Table 11: Baseline average weekly level per ZIP code in the pre-event period.
Variable Control Group Treated Group
Transaction Volume 228.833 221.300
Default 8.714 10.418
@@
Figure 10: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) reports estimates for weekly transaction volume, while Panel (b) shows estimates for weekly default rates. Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
Table 12: Difference-in-Differences results for transaction volume and default rate outcomes at the ZIP code level for payday loan data. The analysis evaluates the impact of hurricane event across treated and control groups over two different weekly windows: from 5 weeks before to 7 weeks after the event ([-5 ; 7]) and from 5 weeks before to 12 weeks after the event ([-5 ; 12]). The coefficient of -4.857∗∗ for transaction volume in the [-5 ; 7] window suggests a significant decrease in transaction volume in the treated ZIP codes, corresponding to approximately 2.2% lower transaction volume relative to pre-event levels. The coefficient for default rate in the same window is -1.088∗∗∗, indicating a significant reduction in the default rate, corresponding to approximately 10.4% lower default rates in treated ZIP codes. The longer window, [-5 ; 12], shows even stronger effects with a coefficient of -8.364∗∗∗ for transaction volume, suggesting a 3.8% reduction in transactions, and -1.458∗∗∗ for default rates, corresponding to a 14% reduction in defaults.
Weekly Window [-5 ; 7] [-5 ; 12]
Dependent Variable Transaction Default Transaction Default
Treated×$\times$×Post -4.857∗∗ -1.088∗∗∗ -8.364∗∗∗ -1.458∗∗∗
[2.040] [0.403] [2.121] [0.416]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Observations 25856 25856 35912 35912
Figure 11: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level for the treated only. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) presents transaction estimates for ZIP codes that received FEMA aid and were in the storm path, while panel Panel (b) shows default estimates for ZIP codes that received FEMA aid and were in the storm path. Standard errors are clustered at the ZIP code level.
Dependent Variable Transaction Default
Period Weekly Weekly
Window [-5 ; 12] [-5 ; 12]
FEMA×$\times$×Post -12.597 -3.316∗∗∗
[7.130] [1.112]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 7366 7366
Table 13: Results for transaction volume and default rate outcomes at the ZIP code level for the payday loan data. The analysis evaluates the impact of hurricane event across treated groups and interacted with receiving FEMA rental assistance over the weekly window: from 5 weeks before to 12 weeks after the event ([-5 ; 12]). The coefficient of -12.597 for transaction volume in the suggests a significant decrease in transaction volume in the treated ZIP codes that received FEMA assistance, corresponding to approximately 5.7% lower transaction volume relative to pre-event levels. The coefficient for default rate in the same window is -3.316∗∗∗, indicating a significant reduction in the default rate, corresponding to approximately 31.8% lower default rates in treated ZIP codes that received FEMA assistance.
@@
Figure 12: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level excluding ZIP codes that received FEMA assistance. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) presents transaction volume estimates for ZIP codes relative to the control group excluding treated ZIP codes that received FEMA assistance, while panel Panel (b) shows default estimates for ZIP codes relative to the control group excluding treated ZIP codes that received FEMA assistance. Standard errors are clustered at the ZIP code level.
Dependent Variable Transaction Default
Period Weekly Weekly
Window [-5 ; 12] [-5 ; 12]
Treated×$\times$×Post -8.538∗∗∗ -0.433
[1.727] [0.348]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 31686 31686
Table 14: Difference-in-differences esults for transaction volume and default rate outcomes at the ZIP code level for the payday loan data. The analysis evaluates the impact of hurricane event relative to the control group excluding treated ZIP codes receiving FEMA rental assistance over the weekly window: from 5 weeks before to 12 weeks after the event ([-5 ; 12]). The coefficient of -8.538 for transaction volume in the suggests a significant decrease in transaction volume in the treated ZIP codes that did not receive FEMA assistance relative to the control group, corresponding to approximately 3.9% lower transaction volume relative to pre-event levels. The coefficient for default rate in the same window is -0.433, indicating a non statistically significant reduction in the default rate when excluding treated ZIP codes that received FEMA assistance.
@@ 3.2.2 Hurricanes and Tropical Storms Combined
Table 15: Baseline average weekly level per ZIP code in the pre-event period for hurricanes and tropical storms combined
Variable Control Group Treated Group
Transaction Volume 267.802 256.220
Default 8.681 9.537
Table 16: Baseline average weekly level per ZIP code in the pre-event period for tropical storms
Variable Control Group Treated Group
Transaction Volume 286.228 291.139
Default 8.666 8.655
Adding lower intensity tropical storms once again does not overturn the results in comparing events that include FEMA assistance. We continue to observe a decrease in transaction volume and a comparative decrease in default rates for events that saw FEMA assistance. Hence, allowing us to control for possible additional temporal events and for differences possibly due to intensity of non-FEMA hurricane events.
@@
Table 17 and Table 18 show that overall transaction volumes decrease by 7.468 weekly transactions per ZIP code and that defaults decrease by 0.982 defaults per week on average relative to a statistically non-significant change in defaults when only including tropical storm events.
Most of the change in decreases in transaction volumes and in preservation of default rates is not attributed to non-FEMA events being possibly due to lower intensity storms relative to FEMA declared hurricane events as tropical storms show similar magnitudes of changes in transaction volumes and non-decrease in default rates to non-FEMA hurricanes as in Table 14.
@@
Figure 13: The figure provides the estimated coefficients from an event study examining the impact of storm (hurricane and tropical storm) exposure on weekly payday loan-related outcomes at the ZIP code level. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) reports estimates for weekly transaction volume, while Panel (b) shows estimates for weekly default rates. Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
Figure 14: The figure provides the estimated coefficients from an event study examining the impact of tropical storm exposure on weekly payday loan-related outcomes at the ZIP code level. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) reports estimates for weekly transaction volume, while Panel (b) shows estimates for weekly default rates. Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
Table 17: Difference-in-Differences results for weekly payday loan transaction volume and default rate outcomes at the ZIP code level, combining both hurricanes and tropical storms. The analysis includes two specifications: one with all storm events (hurricanes and tropical storms), and another excluding hurricane events. The estimation window spans from 5 weeks before to 12 weeks after the event. In the full specification including hurricanes and tropical storms, the coefficient on Treated×$\times$×Post indicates a statistically significant decline of 7.468 transactions and 0.932 defaults in treated ZIP codes. These correspond to approximately 2.9% and 9.8% decreases relative to the treated group’s pre-event averages of 256.220 transactions and 9.537 defaults, respectively. When hurricanes are excluded, the estimated effects are smaller: a 6.345 decrease in transactions (2.2%) and a non-significant 0.273 decrease in defaults (3.2%).
Weekly Window [-5 ; 12] [-5 ; 12]
Dependent Variable Transaction Default Transaction Default
Treated×$\times$×Post -7.468∗∗∗ -0.932∗∗∗ -6.345∗∗∗ -0.273
[1.316] [0.234] [1.847] [0.200]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Hurricanes Y Y N N
Observations 103,291 103,291 67,379 67,379
${}^{*}p<0.1$, ${}^{**}p<0.05$, ${}^{***}p<0.01$
Table 18: Triple Difference-in-Differences results for weekly payday loan transaction volume and default rate outcomes at the ZIP code level, combining both hurricanes and tropical storms. The analysis investigates heterogeneity in the treatment effect by interacting the treatment indicator with a binary storm type indicator (1 = tropical storm, 0 = hurricane). The estimation window spans from 5 weeks before to 12 weeks after the event. The coefficient on Treated×$\times$×Post indicates a statistically significant decrease of 8.364 transactions and 1.458 defaults in treated ZIP codes overall. The interaction term Treated×$\times$×Post×$\times$×Storm captures differential effects by storm type: default rates rise after tropical storms relative to hurricanes, potentially reflecting the role of FEMA assistance, which is less commonly triggered by tropical storms.
Weekly Window [-5 ; 12]
Dependent Variable Transaction Default
Treated×$\times$×Post -8.364∗∗∗ -1.458∗∗∗
[2.040] [0.403]
Treated×$\times$×Post×$\times$×Storm 2.019 1.185∗∗
[2.907] [0.463]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 103,291 103,291
${}^{*}p<0.1$, ${}^{**}p<0.05$, ${}^{***}p<0.01$

### v3

@@ 3.1.2 Hurricanes and Tropical Storms
We expand our baseline results to include tropical storms to show that the increase in observed evictions and filings when lower intensity storms are included no longer holds. We view this as a way to gauge the extent that potential damage severity maybe driving our results, suggesting that the sensitivity to the impact of availability for FEMA assistance matters more within event as opposed to across events. In other words, the relative damage of being on the peripheral distances and demand for assistance has observable impacts on evictions relative to any location for lower intensity storms that do not have FEMA assistance at all. Hence, our result is likely to be robust from other temporal patterns that may have coincided in our analysis with the comparatively fewer hurricane events. Adding additional storm events ensures the increased evictions and filings results are likely due to coinciding with intense hurricane events and the additional absence of FEMA assistance for some geographies.
@@
Figure 9 and Table 7 shows the baseline result when including the greater number of tropical storm events no longer holds (note the three times increase in sample size). Whereas separating the two, shows even a possible statistically significant decline in evictions along the path of tropical storms for radii under 26 nautical miles.
@@
Figure 9: The figure provides the estimated coefficients from an event study examining the combined effect of storm exposure, including hurricanes and tropical storms on monthly eviction-related outcomes at the ZIP code level. The analysis covers a window from five months before to six months after the storm event, with the month immediately preceding the event (month -1) serving as the reference period. Panel (a) displays estimates for monthly eviction rates, while Panel (b) shows estimates for monthly filing rates at the 21 nautical mile treatment radius, whereas Panel (c) displays estimates for monthly eviction rates, while Panel (d) shows estimates for monthly filing rates at the 50 nautical mile treatment radius. Vertical lines indicate 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
Treatment Radius
10 15 21 26 50 100
Hurricanes and Tropical Storms
Panel A: Eviction
[…]
[0.061] [0.059] [0.054] [0.057] [0.051] [0.043]
Panel B: Filing
[…]
[0.105] [0.099] [0.090] [0.079] [0.072] [0.058]
Separate Effects: Hurricane and Tropical Storms
Panel C: Eviction
[…]
[0.100] [0.097] [0.103] [0.114] [0.114] [0.084]
[…]
[0.075] [0.071] [0.060] [0.060] [0.052] [0.048]
Panel D: Filing
[…]
[0.132] [0.126] [0.125] [0.137] [0.150] [0.149]
[…]
[0.114] [0.110] [0.095] [0.078] [0.069] [0.062]
Event-specific ZIP Code FE Y Y Y Y Y Y
Event-specific Time FE Y Y Y Y Y Y
Observations 80,963 83,704 81,921 80,495 75,703 71,206
Table 7: Difference-in-Differences estimates of eviction and filing outcomes following hurricanes and tropical storms. The top panel reports combined treatment effects for all events, while the bottom panels separate effects for hurricanes and tropical storms. Standard errors clustered at the ZIP code level are in brackets. ***, **, and * indicate significance at the 1%, 5%, and 10% levels, respectively.
@@ 3.2.1 Hurricanes
We further compare the usage of high cost credit during hurricane events with the availability of FEMA rental assistance. In particular, as the individual rental assistance provided by FEMA is of similar magnitudes to the average loan amount available to households via payday loans [^8] The maximum amount in the state of Florida is \$500 and often households borrow at the maximum. We find that both transaction volumes and default rates are lower post hurricane event and that this decrease in usage and decrease in defaults is particularly driven when FEMA assistance is made available. Both transactions volumes and defaults decrease by greater magnitudes compared to when it is not made available, reflecting a reduction in demand of payday credit when FEMA is available but also a decrease in precarity in making repayments. In particular, Figure 11 we observe a decrease in defaults and reduction in transaction volumes.
Specifically, in Table 8 we show that the volume of transactions reduce by a maximal amount of 17.609 per week and a decrease in defaults by 2.414 per week on average per ZIP code. This reflects a 8.3% decrease in transactions per week relative to the baseline period and 22.6% decrease in defaults relative to the baseline period.
@@
Comparatively, when FEMA rental assistance is available, we find transactions volumes decrease by 21.209 per week and that non-FEMA ZIP codes see an increase of 13.609 transactions per week at the most maximal, a decline of 10.4% for FEMA receiving ZIP codes and an increase of 6.3% for non-FEMA receiving ZIP codes relative to the pre-event baseline. Furthermore, we find FEMA receiving ZIP codes result in a decrease in defaults by 2.810 per week and that non-FEMA ZIP codes no statistically signficant change across almost all treatment radius. This represents approximately a 26.3% decrease in defaults relative to the pre-event baseline for FEMA receiving ZIP codes.
@@
Figure 10: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level with a treatment radius of 21 nautical miles. The analysis spans a window from five weeks before to twelve weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) reports estimates for weekly transaction volume, while Panel (b) shows estimates for weekly default rates. Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
@@
Figure 11: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level with a treatment radius of 50 nautical miles. The analysis spans a window from five weeks before to twelve weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) reports estimates for weekly transaction volume, while Panel (b) shows estimates for weekly default rates. Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
Treatment Radius
10 15 21 26 50 100
Panel A: Transaction Volume
[…]
[2.962] [2.497] [2.404] [2.228] [2.021] [2.535]
Panel B: Default
[…]
[0.586] [0.479] [0.441] [0.409] [0.404] [0.586]
Event-specific ZIP Code FE Y Y Y Y Y Y
Event-specific Time FE Y Y Y Y Y Y
Observations 45,526 41,821 38,920 36,613 31,904 29,527
Table 8: Difference-in-Differences estimates of post-event impacts on payday loan transaction volume and defaults across treatment radii. Standard errors clustered at the ZIP code level are in brackets. ***, **, and * indicate significance at the 1%, 5%, and 10% levels, respectively.
Treatment Radius
10 15 21 26 50 100
Panel A: Transaction Volume
Control 251.098 259.797 263.990 271.601 291.747 296.754
Treated 247.778 226.742 227.980 217.360 204.819 212.644
Panel B: Default
Control 9.697 9.821 9.713 9.881 10.185 10.085
Treated 12.010 11.146 11.243 10.566 10.298 10.678
Table 9: Baseline averages by group (pre-hurricane period) for each treatment radius. Control and treated ZIP codes are compared across radii of 10, 15, 21, 26, 50, and 100 miles.
Figure 12: Estimated coefficients from an event study model evaluating the impact of hurricane exposure on weekly payday loan transaction volume at the ZIP code level with a treatment radius of 21 nautical miles. The event window spans from five weeks before to twelve weeks after landfall, with the week immediately preceding the hurricane (week -1) serving as the reference period. The treated group is divided based on FEMA rental assistance availability. Standard errors are clustered at the ZIP code level.
Figure 13: Estimated coefficients from an event study model evaluating the impact of hurricane exposure on weekly payday loan transaction volume at the ZIP code level with a treatment radius of 21 nautical miles. The event window spans from five weeks before to twelve weeks after landfall, with the week immediately preceding the hurricane (week -1) serving as the reference period. The treated group is divided based on FEMA rental assistance availability. Standard errors are clustered at the ZIP code level.
Treatment Radius
10 15 21 26 50 100
Panel A: Transaction Volume
[…]
[4.350] [3.096] [2.927] [3.137] [3.096] [3.353]
[…]
[3.412] [3.002] [2.879] [2.525] [2.240] [2.597]
Panel B: Default
[…]
[1.211] [0.809] [0.727] [0.701] [0.632] [0.680]
[…]
[0.654] [0.553] [0.509] [0.458] [0.440] [0.608]
Event-specific ZIP Code FE Y Y Y Y Y Y
Event-specific Time FE Y Y Y Y Y Y
Observations 45,526 41,821 38,920 36,613 31,904 29,527
Table 10: Difference-in-Differences estimates of post-event transaction Volume and default outcomes by FEMA-designated and non-FEMA-designated areas across treatment radii. Standard errors clustered at the ZIP code level are in brackets. ***, **, and * indicate significance at the 1%, 5%, and 10% levels, respectively.
@@ 3.2.2 Hurricanes and Tropical Storms Combined
Adding lower intensity tropical storms for payday loans show a dampened effect for transaction volumes and defaults, unlike the case with evictions where the effect was non-significant. In this case, the results remain significant but the magnitudes decrease with treatment radii, which is in line with the a decrease in potential damage from the storm and the absence of FEMA intervention.
@@
Figure 14 and Table 11 show that overall transaction volumes decrease by a maximal amount of 10.810 weekly transactions per ZIP code and that defaults decrease by 0.504 defaults per week on average.
@@
Figure 14: The figure provides the estimated coefficients from an event study examining the impact of combined hurricane and tropical storm exposure on weekly payday loan-related outcomes at the ZIP code level. The analysis spans a window from five weeks before to twelve weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) reports estimates for weekly transaction volume, while Panel (b) shows estimates for weekly default rates at 21 nautical miles. Panel (c) reports estimates for weekly transaction volume, while Panel (d) shows estimates for weekly default rates at 50 nautical miles. Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
Treatment Radius
10 15 21 26 50 100
Hurricanes and Tropical Storms
Panel A: Transaction Volume
[…]
[1.696] [1.366] [1.262] [1.163] [0.991] [1.285]
Panel B: Default
[…]
[0.304] [0.246] [0.214] [0.187] [0.146] [0.183]
Separate Effects: Hurricane and Tropical Storms
Panel C: Transaction Volume
[…]
[2.962] [2.496] [2.404] [2.228] [2.021] [2.535]
[…]
[1.886] [1.721] [1.519] [1.447] [1.251] [1.468]
Panel D: Default
[…]
[0.585] [0.479] [0.441] [0.409] [0.404] [0.586]
[…]
[0.171] [0.156] [0.140] [0.133] [0.118] [0.176]
Event-specific ZIP Code FE Y Y Y Y Y Y
Event-specific Time FE Y Y Y Y Y Y
Observations 133,020 128,028 131,732 128,489 121,955 116,037
Table 11: Difference-in-Differences estimates of payday loan outcomes following hurricanes and tropical storms. The top panels report combined treatment effects, while the bottom panels separate effects by hurricane and tropical storm exposure. Standard errors clustered at the ZIP code level are in brackets. ***, **, and * indicate significance at the 1%, 5%, and 10% levels, respectively.